The points (a,0) , (0,b) and (1,1) are collinear,If m+n =1, where m and n respectively are- (a) b,-a (b) a,-b (c) 1/a,1/b (d) -1/a,-1/b

Grade:11

1 Answers

Arun
6 years ago
Dear Piyush
 
If these three points are collinear then area of triangle should be zero.
 
Hence after this condition
We get 
1/a + 1/b = 1
Hence when comparing with m+ n = 1
Then
m = 1/a , n = 1/b 
 
Regards
Arun (askIITians forum expert)